EDITORIAL

Terry Goodman

As this editorial is being written, we are in our last week of classes here at Central Missouri State University. With final exams fast approaching, this is a very busy time for us and I am sure it is for you as well. We hope that your busy schedule will not force you to put off reading this issue for very long. Let me now touch on several issues of some importance for MJMS and its readers.

We have received much positive feedback concerning the changes made in the editing and reviewing process for MJMS. Our five associate editors will help us shorten the time from submission to disposition of manuscripts.

Within the next year, we will probably need to increase the subscription rates for the journal. Our rates have remained constant since the beginning of MJMS in 1988. Many of our costs, of course, have increased over that time. We will attempt to keep this rate increase as small as possible.

Many of you have indicated that you enjoy the Problems section of MJMS. Curtis Cooper does an excellent job of editing that section. Let me encourage you to send problems to us over the summer months. Also, you may be able to use the extra (?) time you have this summer to develop solutions for some of the problems appearing in this issue. Finally, please encourage your students to work on the problems as well.

The editorial staff of MJMS wish you a relaxing and productive summer and we hope you look forward to our fall issue.